OpenStudy (anonymous):

how do you do the slope for this y= 1/2 + 2

OpenStudy (anonymous):

Is it 1/2x + 2? Or is it y=1/2 + 2

OpenStudy (anonymous):

If there is no x then your slope is 0, because it is just a horizontal line.

OpenStudy (anonymous):

sorry my bad it's y= 1/2x + 2

OpenStudy (lizz123):

 That is wht is should look first you find the y-intercept which is 2 and then do rise over run for the fraction go up 1 and over 2